# MP1: Mean Flow Tames Policy Learning in 1-step for Robotic Manipulation

### v1

@@ Abstract
In robot manipulation, robot learning has become a prevailing approach. However, generative models within this field face a fundamental trade-off between the slow, iterative sampling of diffusion models and the architectural constraints of faster Flow-based methods, which often rely on explicit consistency losses. To address these limitations, we introduce MP1, which pairs 3D point-cloud inputs with the MeanFlow paradigm to generate action trajectories in one network function evaluation (1-NFE). By directly learning the interval-averaged velocity via the “MeanFlow Identity,” our policy avoids any additional consistency constraints. This formulation eliminates numerical ODE-solver errors during inference, yielding more precise trajectories. MP1 further incorporates CFG for improved trajectory controllability while retaining 1-NFE inference without reintroducing structural constraints. Because subtle scene-context variations are critical for robot learning, especially in few-shot learning, we introduce a lightweight Dispersive Loss that repels state embeddings during training, boosting generalization without slowing inference. We validate our method on the Adroit and Meta-World benchmarks, as well as in real-world scenarios. Experimental results show MP1 achieves superior average task success rates, outperforming DP3 by 10.2% and FlowPolicy by 7.3%. Its average inference time is only 6.8 ms—19× faster than DP3 and nearly 2× faster than FlowPolicy. Our code is available at https://mp1-2254.github.io/.
@@ Introduction
This formulation is key to bypassing iterative integration. Training a network $u_{\theta}$ to model this field directly from its integral definition is intractable. Instead, MeanFlow leverages the “MeanFlow Identity,” a local differential equation derived by differentiating the integral definition to $t$:
@@
While the MeanFlow objective $\mathcal{L}_{cfg}$ excels at learning the temporal dynamics required to produce accurate output trajectories, it provides only an indirect signal for learning the complex mapping from high-dimensional conditional inputs $\mathbf{c}$ to actions. This can lead to a form of ”feature collapse,” where the policy network maps distinct environmental states that demand fundamentally different actions to nearly identical points in its latent space. Such ambiguity is particularly detrimental in robot learning, where subtle differences in object pose or scene configuration are critical for success, especially in few-shot learning regimes.
To directly address this, we incorporate Dispersive Loss, a principled and self-contained regularizer that operates on the model’s internal representations. The core idea is to encourage the latent embeddings of different input samples within a training batch to disperse, thereby enforcing a more discriminative feature space. Dispersive Loss functions as a ”contrastive loss without positive pairs”; the repulsive force is supplied by the loss itself, while the alignment of a state to its correct action is handled implicitly by the primary $\mathcal{L}_{cfg}$ regression objective. Let $\mathbf{z}_{\mathbf{A}}=f_{policy}(\mathbf{A}_{t},t,\mathbf{c})$ be the intermediate representation from a chosen layer within our policy network. The Dispersive Loss is defined as:
@@
Tab. 1 demonstrates that the MP1 achieves SOTA performance across all sub-tasks. The overall average success rate reaches 78.9%$\pm$2.1%, which significantly outperforms the previous best method, FlowPolicy, at 71.6%$\pm$3.5%. Compared to existing approaches, our method yields a 10.2% improvement over DP3 and a 7.3% improvement over FlowPolicy in terms of average success rate, indicating that mean-velocity flows are better suited for the robot learning. In the Meta-World tasks, it can be observed that the MP1 achieves a higher success rate on ”Very Hard” tasks compared to its performance on ”Hard” tasks. The proposed method demonstrates notable improvements across 13 tasks of Meta-World (Medium, Hard, and Very Hard), with success rates increased by 9.8%, 17.9%, and 15.0% respectively over the FlowPolicy. On the 21 ”Easy” tasks in Meta-World, the proposed approach achieves a success rate of 88.2%, representing a 3.4% improvement over the FlowPolicy. Moreover, the proposed method maintains a consistently low standard deviation on certain subtasks (for example, the standard deviation for Adroit Hammer, Door, and the Easy tasks in Meta-World ranges from only 0 to 2.0%), and the average fluctuation in success rate is merely 2.1%, which is lower than that of other methods. These results further indicate that the proposed approach exhibits greater stability and reliability under different random seeds.

### v2

@@ Abstract
In robot manipulation, robot learning has become a prevailing approach. However, generative models within this field face a fundamental trade-off between the slow, iterative sampling of diffusion models and the architectural constraints of faster Flow-based methods, which often rely on explicit consistency losses. To address these limitations, we introduce MP1, which pairs 3D point-cloud inputs with the MeanFlow paradigm to generate action trajectories in one network function evaluation (1-NFE). By directly learning the interval-averaged velocity via the “MeanFlow Identity”, our policy avoids any additional consistency constraints. This formulation eliminates numerical ODE-solver errors during inference, yielding more precise trajectories. MP1 further incorporates CFG for improved trajectory controllability while retaining 1-NFE inference without reintroducing structural constraints. Because subtle scene-context variations are critical for robot learning, especially in few-shot learning, we introduce a lightweight Dispersive Loss that repels state embeddings during training, boosting generalization without slowing inference. We validate our method on the Adroit and Meta-World benchmarks, as well as in real-world scenarios. Experimental results show MP1 achieves superior average task success rates, outperforming DP3 by 10.2% and FlowPolicy by 7.3%. Its average inference time is only 6.8 ms—19× faster than DP3 and nearly 2× faster than FlowPolicy. Our code is available at https://github.com/LogSSim/MP1.git.
@@ Introduction
This formulation is key to bypassing iterative integration. Training a network $u_{\theta}$ to model this field directly from its integral definition is intractable. Instead, MeanFlow leverages the “MeanFlow Identity”, a local differential equation derived by differentiating the integral definition to $t$:
@@
While the MeanFlow objective $\mathcal{L}_{cfg}$ excels at learning the temporal dynamics required to produce accurate output trajectories, it provides only an indirect signal for learning the complex mapping from high-dimensional conditional inputs $\mathbf{c}$ to actions. This can lead to a form of “feature collapse”, where the policy network maps distinct environmental states that demand fundamentally different actions to nearly identical points in its latent space. Such ambiguity is particularly detrimental in robot learning, where subtle differences in object pose or scene configuration are critical for success, especially in few-shot learning regimes.
To directly address this, we incorporate Dispersive Loss, a principled and self-contained regularizer that operates on the model’s internal representations. The core idea is to encourage the latent embeddings of different input samples within a training batch to disperse, thereby enforcing a more discriminative feature space. Dispersive Loss functions as a “contrastive loss without positive pairs”; the repulsive force is supplied by the loss itself, while the alignment of a state to its correct action is handled implicitly by the primary $\mathcal{L}_{cfg}$ regression objective. Let $\mathbf{z}_{\mathbf{A}}=f_{policy}(\mathbf{A}_{t},t,\mathbf{c})$ be the intermediate representation from a chosen layer within our policy network. The Dispersive Loss is defined as:
@@
Tab. 1 demonstrates that the MP1 achieves SOTA performance across all sub-tasks. The overall average success rate reaches 78.9%$\pm$2.1%, which significantly outperforms the previous best method, FlowPolicy, at 71.6%$\pm$3.5%. Compared to existing approaches, our method yields a 10.2% improvement over DP3 and a 7.3% improvement over FlowPolicy in terms of average success rate, indicating that mean-velocity flows are better suited for the robot learning. In the Meta-World tasks, it can be observed that the MP1 achieves a higher success rate on “Very Hard” tasks compared to its performance on “Hard” tasks. The proposed method demonstrates notable improvements across 13 tasks of Meta-World (Medium, Hard, and Very Hard), with success rates increased by 9.8%, 17.9%, and 15.0% respectively over the FlowPolicy. On the 21 “Easy” tasks in Meta-World, the proposed approach achieves a success rate of 88.2%, representing a 3.4% improvement over the FlowPolicy. Moreover, the proposed method maintains a consistently low standard deviation on certain subtasks (for example, the standard deviation for Adroit Hammer, Door, and the Easy tasks in Meta-World ranges from only 0 to 2.0%), and the average fluctuation in success rate is merely 2.1%, which is lower than that of other methods. These results further indicate that the proposed approach exhibits greater stability and reliability under different random seeds.